Find all the points having a y-coordinate of -6 whose distance is from the point (1,2) is 17
Answer by ewatrrr(24785)

 
Hi,
(1 ,2)
(x,-6) (1-x)^2 + 8^2 = 17^2
1 - 2x + x^2 = 225
x^2 - 2x - 224 = 0
(x - 16)(x+14) = 0
x = 16 0r x = -14
Find all the points having a y-coordinate of -6 whose distance is from the point (1,2) is 17
Pt(16,-6) and Pt(-14,-6)